OpenStudy (anonymous):

Solve each equation for x by converting to exponential form. In part (b), give two forms for the answer: one involving e and the other a calculator approximation rounded to two decimal places. (a) log5(x) = −2 (b) ln(x) = −2

OpenStudy (anonymous):

(a)= 1/25 b) = 1/e^2

OpenStudy (anonymous):

(a)\[\log_5(x) = -2 \Rightarrow 5^{\log_5(x)} = 5^{-2} \Rightarrow x = \frac{1}{25}\](b) \[\log(x) = -2 \Rightarrow e^{\log(x)} = e^{-2} = \frac{1}{e^2} \approx 0.14.\]

OpenStudy (dumbcow):

In general use this to convert from log to exponential or vis versa \[\log_{b}x = y \leftarrow \rightarrow x = b^{y} \]